# Global ${\mathcal{L}}^{2}$ minimization at uniform exponential rate via geometrically adapted gradient descent in Deep Learning

### v4

@@ 1. Introduction and Main Results
Our main goal here is to develop a rigorous mathematical understanding of some crucial dynamical systems aspects of supervised learning in DL networks; we do not focus on the development of efficient algorithms, or exploration of specific application cases. The definition of a gradient vector field necessitates the choice of a Riemannian metric structure, a general fact of differential geometry. In the standard approach to DL, the gradient of the cost function is determined using the Euclidean metric on the space of parameters (weights and biases); this choice is ad hoc, and not motivated by geometric reasons. As explained in (1.25), below, this defines a gradient flow in the output layer of the DL network with respect to a highly nontrivial induced metric; the latter is the reason for the trapping of orbits.
@@
In the overparametrized case, we prove in Theorem 2.1 that, provided a rank condition holds, all orbits of the modified gradient flow drive the ${\mathcal{L}}^{2}$ cost to its global minimum at a uniform exponential convergence rate; this yields an a priori stopping time for any prescribed proximity to the global ${\mathcal{L}}^{2}$ cost minimum. In Section 2.4, we point out relations of the latter to sub-Riemannian geometry. In Section 3, we address the trapping of orbits by approximate local minima in the standard gradient descent flow. In Section 4, we generalize the above framework to the situation in which the rank condition does not hold; in particular, we show that local equilibria can only exist if a rank loss occurs, and that generically, they are not isolated points, but elements of a critical submanifold of parameter space. In the underparametrized situation, we show in Theorem 5.1 that the analogous version of the modified gradient flow is mapped to a geometrically natural constrained gradient flow. In Section 6, we show that in the borderline case separating the strictly overparametrized and underparametrized situations, the two modified gradient flows coincide. A key objective in this work is to elucidate and exploit the link between gradient flows in the parameter space and in the output layer.
@@ 1.1. Definition of the mathematical model
via the weight matrix $W_{\ell}\in{\mathbb{R}}^{M_{\ell}\times M_{\ell-1}}$, bias vector $b_{\ell}\in{\mathbb{R}}^{M_{\ell}}$, and activation function $\sigma$. We assume that $\sigma$ has a Lipschitz continuous derivative, and that the output layer
@@
contains no activation function.
@@
the cost ${\mathcal{C}}[{\underline{x}}[\underline{\theta}(s)]]$ is monotone decreasing in $s\in{\mathbb{R}}_{+}$, and since ${\mathcal{C}}[{\underline{x}}[\underline{\theta}(s)]]\geq 0$ is bounded below, the limit ${\mathcal{C}}_{*}=\lim_{s\rightarrow\infty}{\mathcal{C}}[{\underline{x}}[\underline{\theta}(s)]]$ exists for any orbit $\{\underline{\theta}(s)|s\in{\mathbb{R}}\}$, and depends on the initial data, ${\mathcal{C}}_{*}={\mathcal{C}}_{*}[{\underline{x}}[\underline{\theta}(0)]]$. In applications, the initial data $\underline{\theta}_{0}\in{\mathbb{R}}^{K}$ is often chosen at random.
@@
with initial data $x_{j}(0)=x_{j}^{(0)}$, so that using the expression (1.7), we find
@@ 1.4. Geometrically adapted gradient descent
The right hand side of (1.25) is the gradient vector field defined with respect to the Riemannian structure on ${\mathbb{R}}^{QN}$ with metric tensor $(DD^{T})^{-1}$. This is the pushforward metric on ${\mathbb{R}}^{QN}$ induced by the Euclidean metric on ${\mathbb{R}}^{K}$ associated to (1.22), induced by the map ${\mathbb{R}}^{K}\rightarrow{\mathbb{R}}^{QN}$, $\underline{\theta}\mapsto{\underline{x}}[\underline{\theta}]$.
We will discuss this fact in more detail in Section 3. This induced metric is the geometric origin of the complicated cost landscape; orbits can get trapped in approximate local minima (locations where the norm of the gradient is small) which typically occur in large quantities, and whose distribution is very difficult to control.
@@ 1.4.1. The overparametrized case
We note that the right hand side of (1.29) is in the range of $P[\underline{\theta}(s)]$, and therefore has the structure of a constrained dynamical system in ${\mathbb{R}}^{K}$.
@@
The geometric structure will be described in detail in Section 2; the map ${\underline{x}}:{\mathbb{R}}^{K}\rightarrow{\mathbb{R}}^{QN}$ defines the pullback vector bundle ${\mathcal{V}}\subset T{\mathbb{R}}^{K}$, endowed with the pullback metric $h$ induced by the Euclidean structure on ${\mathbb{R}}^{QN}$. Then, the operator ${\rm Pen}[D]({\rm Pen}[D)^{T}\nabla_{\underline{\theta}}$ on the right hand side of (1.29) is the coordinate representation of the gradient ${\rm grad}_{h}$ on ${\mathcal{V}}$. This ensures that the vector field (1.29) is horizontal with respect to ${\mathcal{V}}$, and in particular, that its pushforward under ${\underline{x}}$ is the gradient vector field $-\nabla_{{\underline{x}}}{\mathcal{C}}[{\underline{x}}]$ on $T{\mathbb{R}}^{QN}$ with the Euclidean metric. As we will explain in Section 2.4, ${\mathcal{V}}$ is in general non-holonomic, and the triple $({\mathbb{R}}^{K},{\mathcal{V}},h)$ defines a sub-Riemannian manifold.
@@ 2. Geometrically adapted gradient flow for overparametrized DL
Thus, it follows from the discussion in Section 1.3 that if (2.1) is satisfied along the orbit $\underline{\theta}(s)$ for all $s\in{\mathbb{R}}$, then
@@
Convexity of ${\mathcal{C}}[{\underline{x}}]$ in ${\underline{x}}-{\underline{y}}_{\omega}$ implies that $\lim_{s\rightarrow}{\underline{x}}[\underline{\theta}(s)]={\underline{y}}_{\omega}$ converges. The convergence rates are as in (1.13) and (1.14), because (2.9) is the same dynamical system as (1.11).
@@ 2.1. Rank condition on $DD^{T}$
The rank of $DD^{T}$ plays a major role in the above considerations. Conjecturally, one may expect that with high probability, $DD^{T}$ exhibits full rank (that is, ${\rm rank}(DD^{T})=QN$), when $K>QN$ is sufficiently large. To see this, we note that
$D[\underline{\theta}]D^{T}[\underline{\theta}]$ $=$ $\sum_{j=1}^{K}\partial_{\theta_{j}}{\underline{x}}[\underline{\theta}](\partial_{\theta_{j}}{\underline{x}}[\underline{\theta}])^{T}$ (2.12)
$=$ $\sum_{j=1}^{K}|\partial_{\theta_{j}}{\underline{x}}[\underline{\theta}]|_{{\mathbb{R}}^{QN}}^{2}e_{j}[\underline{\theta}]e_{j}^{T}[\underline{\theta}]$
where $e_{j}[\underline{\theta}]e_{j}^{T}[\underline{\theta}]$ is the rank one projection onto the span of $e_{j}[\underline{\theta}]:=\frac{\partial_{\theta_{j}}{\underline{x}}[\underline{\theta}]}{|\partial_{\theta_{j}}{\underline{x}}[\underline{\theta}|}\in S^{QN-1}\subset{\mathbb{R}}^{QN}$ where $S^{k}$ is the $k$-dimensional unit sphere. The question then is how the random sequence $\{e_{j}[\underline{\theta}]\}_{j=1}^{K}$ is distributed on $S^{QN-1}$.
As a simple example, let us consider the following idealized case.
Proposition 2.2.
Assume that for (almost) every $\underline{\theta}\in{\mathbb{R}}^{K}$, with $K$ sufficiently large, the family of vectors $\{e_{j}[\underline{\theta}]\}_{j=1}^{K}$ is i.i.d. distributed on $S^{QN-1}$ with respect to an absolutely continuous probability measure $\mu_{\underline{\theta}}$. Let the $QN-1$ dimensional open set $U:={\rm supp}(\mu_{\underline{\theta}})\subset S^{QN-1}$ denote its support. Then,
$\mathbb{P}\Big{[}{\rm rank}(D[\underline{\theta}]D^{T}[\underline{\theta}])=QN\Big{]}=1\;\;,\;\;{\rm when}\;K>QN\,.$ (2.13)
Proof.
We argue by induction to prove
$\mathbb{P}\Big{[}{\rm span}\big{(}\{e_{j}[\underline{\theta}]\}_{j=1}^{K}\big{)}={\mathbb{R}}^{QN}\Big{]}=1\;\;,\;\;{\rm when}\;K>QN\,,$ (2.14)
which is equivalent to the claim. Assume for the base case that given a random point $\widetilde{e}_{1}\in U\subset S^{QN-1}$ where $\widetilde{e}_{i}:=e_{j_{i}}$, we pick another random point $\widetilde{e}_{2}\in U$. The probability of $\widetilde{e}_{2}$ being either $\widetilde{e}_{1}$ or $-\widetilde{e}_{1}$ (if the latter is contained in $U$) is
[…]
because the set $\{\widetilde{e}_{1},-\widetilde{e}_{1}\}$ has dimension zero, and thus has measure zero with respect to $\mu_{\underline{\theta}}$. Hence, almost surely, $\widetilde{e}_{1}$ and $\widetilde{e}_{2}$ are linearly independent in ${\mathbb{R}}^{QN}$. Next, assume by induction that we have picked $k\leq QN-1$ points $\{\widetilde{e}_{1},\dots,\widetilde{e}_{k}\}\subset S^{QN-1}$ at random, and that they are linearly independent in ${\mathbb{R}}^{QN}$. Let $H_{k}:={\rm span}(\{\widetilde{e}_{1},\dots,\widetilde{e}_{k}\})\subset{\mathbb{R}}^{QN}$ be the $k$-dimensional plane spanned by them. Then, the probability of picking the $k+1$-st point $\widetilde{e}_{k+1}$ at random so that it is contained in $H_{k}\cap S^{QN-1}$ is
[…]
because $H_{k}\cap U$ is a zero measure set with respect to $\mu_{\underline{\theta}}$, of dimension $k-1<QN-1$. In the terminal step of the induction, we have arrived at $k=QN-1$ random points spanning the hyperplane $H_{QN-1}$. The probability of picking the $QN$-th random point $\widetilde{e}_{QN}\in U$ in the intersection $H_{QN-1}\cap S^{QN-1}$ is zero because its dimension is $QN-2<QN-1$, and therefore is a measure zero set.
Therefore, we conclude from induction that with probability $1$, $K\geq QN$ points picked at random on $U\subset S^{QN-1}$ are linearly independent. ∎
It remains to be investigated to which extent the general case (2.12) exhibits randomness, and whether this simple example has relevance for it.
@@ 3.1. Trapping of orbits
This pushforward metric is the source of the complicated cost landscape encountered for the standard gradient flow, and the emergence of a multitude of approximate local minima (where the norm of $DD^{T}\nabla_{{\underline{x}}}{\mathcal{C}}$ in (3.1) is small) which can trap its orbits. The main issue here arises from the fact that a priori, there exists no uniform lower bound on $DD^{T}$, that is,
@@
from being arbitrarily small, even if $DD^{T}>0$.
As a consequence, there may exist regions that ”trap” orbits of (3.7) for a long time in the following sense, even if ${\rm rank}(D^{T})=QN$ is maximal, so that $DD^{T}>0$ holds.
@@
We write the standard gradient descent flow (1.9) in the form
$\underline{\theta}(0)$ $=$ $\underline{\theta}_{0}\;\in{\mathbb{R}}^{K}$
using (1.21). Clearly,
@@ 3.2. Local extrema and rank loss
Therefore, the existence of an isolated local equilibrium point $\underline{\theta}_{*}$ of (3.7) away from the global minimimum of the cost must satisfy both ${\rm rank}(D[\underline{\theta}_{*}])<QN$ and (3.11). ∎
@@ 4.1.2. Rank loss in the geometrically adapted gradient flow
The definition of the geometrically adapted gradient flow given in Theorem 2.1 needs to be modified in case a rank loss, ${\rm rank}(D)<QN$, is admitted. In fact, we instead define the following differential-algebraic system; for background on the efficient computational simulation of differential-algebraic systems, see for instance [12].
@@
$\partial_{s}\underline{\theta}(s)$ $=$ $-D^{T}[\underline{\theta}(s)]\Psi[\underline{\theta}(s)]$ (4.5)
@@
Then, if $rank(D)=QN$ is maximal, (4.5) is equivalent to the geometrically adapted gradient flow in in Theorem 2.1.
@@
$=$ $-(DD^{T})[\underline{\theta}(s)]\Psi[\underline{\theta}(s)]$
@@
A point $\underline{\theta}_{*}\in{\mathbb{R}}^{K}$ is an equilibrium of the standard gradient flow (3.7) if and only if it is an equilibrium of the geometrically adapted gradient flow (4.5) (respectively (2.1) in case $D$ has full rank).
@@
If a rank loss occurs so that ${\rm rank}(D^{T})<QN$ in a neighborhood $U^{\prime}\subset{\mathbb{R}}^{K}$, then an equilibrium of (4.4) and an equilibrium of (4.7) both require the same condition
@@
For (4.4), this is because of ${\mathcal{P}}DD^{T}=DD^{T}{\mathcal{P}}$, and $DD^{T}$ has full rank on the range of ${\mathcal{P}}$. Therefore, any equilibrium of (2.1) that is not a global minimizer satisfies (3.10).
Next, assume that ${\rm rank}({\mathcal{P}})=r<QN$ on an open neighborhood $U\subset{\mathbb{R}}^{K}$, and let $\underline{\theta}\in U$. Given that the activation function $\sigma$ is smooth, both $D[\underline{\theta}]$ and ${\underline{x}}[\underline{\theta}]$ depend smoothly on $\underline{\theta}$. Assume that the linearly independent family of smooth vector fields $V_{\alpha}:{\mathbb{R}}^{K}\rightarrow{\mathbb{R}}^{QN}$, $\alpha=1,\dots,r$, spans the range of ${\mathcal{P}}$; for instance, one could choose a suitable $r$-tuple of column vectors of $D$. Then, we define the smooth family of functions $g_{\alpha}:{\mathbb{R}}^{K}\rightarrow{\mathbb{R}}$ defined by the inner product
@@
We conclude that the standard gradient flow (3.7) and the geometrically adapted gradient flow (2.1) have the same set of equilibria. If those equilibria are not global, then they are generically not isolated points, but elements of a critical manifold. On the other hand, (2.1) exhibits a uniform exponential convergence rate whenever $DD^{T}$ has full rank, as opposed to (3.7) which provides no a priori control over convergence rates.

### v5

@@ 1. Introduction and Main Results
Our main goal here is to develop a rigorous mathematical understanding of some crucial dynamical systems aspects of supervised learning in DL networks; we do not focus on the development of efficient algorithms, or exploration of specific application cases. The definition of a gradient vector field necessitates the choice of a Riemannian metric structure, a general fact of differential geometry. In the standard approach to DL, the gradient of the cost function is determined using the Euclidean metric on the space of parameters (weights and biases); this choice is ad hoc, and not motivated by geometric reasons. As explained in (1.25), below, this defines a gradient flow in the output layer of the DL network with respect to a highly nontrivial induced metric determined by the Neural Tangent Kernel (NTK), [15]; the latter is the reason for the trapping of orbits.
@@
In the overparametrized case, we prove in Theorem 2.1 that, provided a rank condition holds, all orbits of the modified gradient flow drive the ${\mathcal{L}}^{2}$ cost to its global minimum at a uniform exponential convergence rate; this yields an a priori stopping time for any prescribed proximity to the global ${\mathcal{L}}^{2}$ cost minimum. In Section 2.4, we point out relations of the latter to sub-Riemannian geometry. It is known that invertibility of the NTK implies that zero loss is reachable, [17, 19]. However, trapping of orbits by approximate local minima of the standard gradient descent flow can occur; this is addressed in Section 3. In Section 4, we generalize the above framework to the situation in which the rank condition does not hold; in particular, we show that local equilibria can only exist if a rank loss occurs, and that generically, they are not isolated points, but elements of a critical submanifold of parameter space; related results for global minima of overparametrized networks were established in [9]. In the underparametrized situation, we show in Theorem 5.1 that the analogous version of the modified gradient flow is mapped to a geometrically natural constrained gradient flow. In Section 6, we show that in the borderline case separating the strictly overparametrized and underparametrized situations, the two modified gradient flows coincide. A key objective in this work is to elucidate and exploit the link between gradient flows in the parameter space and in the output layer.
@@ 1.1. Definition of the mathematical model
via the weight matrix $W_{\ell}\in{\mathbb{R}}^{M_{\ell}\times M_{\ell-1}}$, bias vector $b_{\ell}\in{\mathbb{R}}^{M_{\ell}}$, and activation function $\sigma$. We assume that the output layer
@@
contains no activation function, and that $\sigma$ has a Lipschitz continuous derivative (non-smooth activations like ReLU can be accommodated by a local mollification).
@@
the cost ${\mathcal{C}}[{\underline{x}}[\underline{\theta}(s)]]$ is monotone decreasing in $s\in{\mathbb{R}}_{+}$, and since ${\mathcal{C}}[{\underline{x}}[\underline{\theta}(s)]]\geq 0$ is bounded below, the limit ${\mathcal{C}}_{*}=\lim_{s\rightarrow\infty}{\mathcal{C}}[{\underline{x}}[\underline{\theta}(s)]]$ exists for any orbit $\{\underline{\theta}(s)|s\in{\mathbb{R}}_{+}\}$, and depends on the initial data, ${\mathcal{C}}_{0}={\mathcal{C}}[{\underline{x}}[\underline{\theta}(0)]]$. In applications, the initial data $\underline{\theta}_{0}\in{\mathbb{R}}^{K}$ is often chosen at random.
@@
with initial data $x_{j}(0)\in{\mathbb{R}}^{QN}$, so that using the expression (1.7), we find
@@ 1.4. Geometrically adapted gradient descent
The right hand side of (1.25) is the gradient vector field defined with respect to the Riemannian structure on ${\mathbb{R}}^{QN}$ with metric tensor $(DD^{T})^{-1}$. This is the pushforward metric on ${\mathbb{R}}^{QN}$ induced by the Euclidean metric on ${\mathbb{R}}^{K}$ associated to (1.22), induced by the map ${\mathbb{R}}^{K}\rightarrow{\mathbb{R}}^{QN}$, $\underline{\theta}\mapsto{\underline{x}}[\underline{\theta}]$. The matrix $DD^{T}$ corresponds to the Neural Tangent Kernel (NTK), [15].
We will discuss this fact in more detail in Section 3. This induced metric is the geometric origin of the complicated cost landscape; it is known that if $DD^{T}$ is invertible, zero loss can be reached, [17, 19]. However, orbits can get trapped in approximate local minima (locations where the norm of the gradient is small) which typically occur in large quantities, and whose distribution is very difficult to control.
@@ 1.4.1. The overparametrized case
We note that $P\,{\rm Pen}[D]={\rm Pen}[D]D{\rm Pen}[D]={\rm Pen}[D]$ due to (1.27). Therefore, the right hand side of (1.29) is mapped by $P[\underline{\theta}(s)]$ to itself, and lies in the range of $P[\underline{\theta}(s)]$. Hence, (1.29) has the structure of a constrained dynamical system in ${\mathbb{R}}^{K}$.
@@
The geometric structure will be described in detail in Section 2; the map ${\underline{x}}:{\mathbb{R}}^{K}\rightarrow{\mathbb{R}}^{QN}$ defines the pullback vector bundle ${\mathcal{V}}\subset T{\mathbb{R}}^{K}$, endowed with the pullback metric $h$ induced by the Euclidean structure on ${\mathbb{R}}^{QN}$. Then, the operator ${\rm Pen}[D]({\rm Pen}[D])^{T}\nabla_{\underline{\theta}}$ on the right hand side of (1.29) is the coordinate representation of the gradient ${\rm grad}_{h}$ on ${\mathcal{V}}$. This ensures that the vector field (1.29) is horizontal with respect to ${\mathcal{V}}$, and in particular, that its pushforward under ${\underline{x}}$ is the gradient vector field $-\nabla_{{\underline{x}}}{\mathcal{C}}[{\underline{x}}]$ on $T{\mathbb{R}}^{QN}$ with the Euclidean metric. As we will explain in Section 2.4, ${\mathcal{V}}$ is in general non-holonomic, and the triple $({\mathbb{R}}^{K},{\mathcal{V}},h)$ defines a sub-Riemannian manifold.
@@ 2. Geometrically adapted gradient flow for overparametrized DL
Thus, it follows from the discussion in Section 1.3 that if (2.1) is satisfied along the orbit $\underline{\theta}(s)$ for all $s\in{\mathbb{R}}_{+}$, then
@@
Convexity of ${\mathcal{C}}[{\underline{x}}]$ in ${\underline{x}}-{\underline{y}}_{\omega}$ implies that $\lim_{s\rightarrow\infty}{\underline{x}}[\underline{\theta}(s)]={\underline{y}}_{\omega}$ converges. The convergence rates are as in (1.13) and (1.14), because (2.9) is the same dynamical system as (1.11).
@@ 2.1. Rank condition on $DD^{T}$
The rank of $DD^{T}$ plays a major role in the above considerations. In [15], a detailed discussion is provided about the fact that in the infinite width limit of a deep network, the NTK $DD^{T}$ is positive definite, and thus of full rank. Without the assumption of the infinite width limit, it is a subtle and challenging problem to determine sufficient conditions for full rank. In [8], the NTK is proven to have full rank on an open subset of parameter space ${\mathbb{R}}^{K}$ in overparametrized deep networks with sufficient width and for generic training data; but the question whether this property is preserved along the entire gradient descent trajectory remains to be investigated in future work.
@@ 3.1. Trapping of orbits
This pushforward metric is the source of the complicated cost landscape encountered for the standard gradient flow, and the emergence of a multitude of approximate local minima (where the norm of $DD^{T}\nabla_{{\underline{x}}}{\mathcal{C}}$ in (3.1) is small) which can trap its orbits.
We write the standard gradient descent flow (1.9) in the form
$\underline{\theta}(0)$ $=$ $\underline{\theta}_{0}\;\in{\mathbb{R}}^{K}$
using (1.21).
It is known that invertibility of $DD^{T}$ implies that zero loss can be reached, [17, 19]. However, the main issue here arises from the fact that a priori, there exists no uniform lower bound on the NTK $DD^{T}$, that is,
@@
from being arbitrarily small, even if $DD^{T}>0$. As a consequence, there may exist regions that ”trap” orbits of (3.2) for a long time in the following sense, even if ${\rm rank}(D^{T})=QN$ is maximal, so that $DD^{T}>0$ holds.
@@
Clearly,
@@ 3.2. Local extrema and rank loss
Therefore, existence of an isolated local equilibrium point $\underline{\theta}_{*}$ of (3.2), which is not a global minimimum, requires that both ${\rm rank}(D[\underline{\theta}_{*}])<QN$ and (3.11) hold.
@@ 4.1.2. Rank loss in the geometrically adapted gradient flow
The definition of the geometrically adapted gradient flow given in Theorem 2.1 needs to be modified in case a rank loss, ${\rm rank}(D)<QN$, is admitted. In fact, we instead define the following differential-algebraic system; for background on the efficient computational solution of differential-algebraic systems, see for instance [13].
@@
$\partial_{s}\underline{\theta}(s)$ $=$ $D^{T}[\underline{\theta}(s)]\Psi[\underline{\theta}(s)]$ (4.5)
@@
Then, if $rank(D)=QN$ is maximal, (4.5) is equivalent to the geometrically adapted gradient flow in Theorem 2.1.
@@
$=$ $(DD^{T})[\underline{\theta}(s)]\Psi[\underline{\theta}(s)]$
@@
A point $\underline{\theta}_{*}\in{\mathbb{R}}^{K}$ is an equilibrium of the standard gradient flow (3.2) if and only if it is an equilibrium either of the geometrically adapted gradient flow (4.5) if $D$ does not have full rank, or of (2.1) if $D$ has full rank.
@@
If a rank loss occurs, ${\rm rank}(D^{T})<QN$, in a neighborhood $U^{\prime}\subset{\mathbb{R}}^{K}$, then an equilibrium of (4.4) and an equilibrium of (4.7) both require the same condition
@@
For (4.4), this is because of $DD^{T}={\mathcal{P}}DD^{T}{\mathcal{P}}$ has full rank on the range of ${\mathcal{P}}$. Therefore, any equilibrium of (2.1) that is not a global minimizer satisfies (3.10).
Next, assume that ${\rm rank}({\mathcal{P}})=r<QN$ on an open neighborhood $U\subset{\mathbb{R}}^{K}$, and let $\underline{\theta}\in U$. Given that the activation function $\sigma$ is smooth, both $D[\underline{\theta}]$ and ${\underline{x}}[\underline{\theta}]$ depend smoothly on $\underline{\theta}$. Assume that the linearly independent family of smooth vector fields $V_{\alpha}:{\mathbb{R}}^{K}\rightarrow{\mathbb{R}}^{QN}$, $\alpha=1,\dots,r$, spans the range of ${\mathcal{P}}$; for instance, one could choose a suitable $r$-tuple of column vectors of $D$. Then, we define the smooth family of functions $g_{\alpha}:{\mathbb{R}}^{K}\rightarrow{\mathbb{R}}$ obtained from the inner product
@@
We conclude that the standard gradient flow (3.2) and the geometrically adapted gradient flow (2.1) have the same set of equilibria. If those equilibria are not global, then they are generically not isolated points, but elements of a critical manifold. On the other hand, (2.1) exhibits a uniform exponential convergence rate whenever $DD^{T}$ has full rank, as opposed to (3.2) which provides no such a priori control. We also note that in [9], it is shown that global minima in overparametrized networks define critical submanifolds.